Grade 12Trigonometry

If a>0,bx->0+ (root(1-cos 2ax))/sin bx is equal to _______
-a (root2)/b
|a|(root2)/|b|
-|a|(root2)/|b|
PLEASE EXPLAIN THE SOLUTION OF YOUR ANSWER

Aditya Mitra
7 Years agoGrade 12

1 Answer

Aditya Gupta
ApprovedApproved Tutor Answer7 Years ago
root(1-cos 2ax)= root2*sinax
limit is root2*a/b because limx->0+sinax/sinbx=a/b*limx->0+sinax/ax*bx/sinbx=a/b*1*1=a/b
so, answer is root2*a/b